OpenStudy (anonymous):

Your math teacher has a total 8 bonus points to give out to his 2 classes. Based on their effort in the review game, class 1 gets 4 less than 2 times the amount of class 2. How many bonus points does each class get? In order to graph, let class 1 be x and class 2 be y.

OpenStudy (anonymous):

Let class 1 = x and class 2 = y doh it already says that anyways lets dissect this sentence Class 1 gets 4 less than 2 times the amount of class 2 now first "Class 2 is multiplied by 2" so we know we'll have 2y in this equation

OpenStudy (anonymous):

next lets look at "class 1 gets 4 less than 2y so lets take some values.... if 2y = 10, than the amount class 2 gets is 6 because that is 4 less than 2y so to write this down as an equation x will always equal 2 times y minus 4 x=2y-4

OpenStudy (anonymous):

to graph this it'd be easier to solve for y \[x+4=2y\] \[\frac{1}{2}x+2=y\]

OpenStudy (anonymous):

now we have to do the first sentence also.. "your teacher can give out only 8 points to both classes, together".... this means that class 1's points and class 2 points must equal 8 or x+y=8

OpenStudy (anonymous):

solving this for y, you get y=-x +8

OpenStudy (anonymous):

point intercept form is \[y=mx+b\] which is what i did... it's much easier to graph it this way because \[b=y intercept=(0,y)\]

OpenStudy (anonymous):

so can you try graphing this using (0, b) as one point and using rise/run?

OpenStudy (anonymous):

yep both are in intercept form now \[y=\frac{1}{2}x+2\] \[y=-x+8\] can you graph these or is that the step you're having trouble with?

OpenStudy (anonymous):

\[\frac{1}{2}x+2=-x+8\] \[\frac{3}{2}x=6\] \[x=6*\frac{2}{3}\] \[\frac{12}{3}=4\] \[x=4\] using either equation \[y=-x+8\] \[y=-4+8\] \[y=4\] solution is (4,4) you're right =]

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

If you need any more help, let me know